AB passes through the points (1, 3) and (3, 17). If CD is perpendicular to AB, what is the slope of CD?
Answer by lwsshak3(6505)

AB passes through the points (1, 3) and (3, 17). If CD is perpendicular to AB, what is the slope of CD?
**
The slope of CD is the negative reciprocal of the slope of AB.
Slope of AB=∆y/∆x=(17-3)/(3-1)=14/2=7
Slope of CD=-1/7